OpenStudy (nuccioreggie):

The elevation of city A is 96 feet below sea level; the elevation of city C is 24 feet below sea level. The elevation of city D is 1 over 4 of the elevation of city C. Part A: Write an expression to represent the elevation, in feet, of city D. (5 points) Part B: Solve the expression and explain why the answer is negative or positive. (5 points)

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (michele_laino):

part A: the requested elevation is: \[\frac{1}{4} \times 24 = ...?\] @nuccioreggie

OpenStudy (nuccioreggie):

Sorry idont know that

OpenStudy (nuccioreggie):

Well i dont know how to times a fraction

OpenStudy (michele_laino):

what is 24 divided by 4?

OpenStudy (nuccioreggie):

6

OpenStudy (michele_laino):

that's right!

OpenStudy (nuccioreggie):

Ok so that the answer for 1 over 4 × 24 ???

OpenStudy (michele_laino):

for part B, we have to add the right sign close to 6, namely the right answer is -6, since all elevations below the sea level have to be considered negative, so we have: elevation of A = -96 elevation of C = -24 and elevation of D = (1/4)*(-24)= -6 whereas all elevetions above the sea level have to be considered positive

OpenStudy (nuccioreggie):

Ok so all thats the answer for b ?

OpenStudy (michele_laino):

yes!